# Flows of linear orders on sparse graphs

### v1

@@ 4.3. Proof idea - informal overview
The incompatibility of the two ordered graphs $C_{0}^{\prec},C_{1}^{\prec}$ in $\mathcal{J}$ will result from them forcing incompatible orientations: we can use the order $\prec_{\beta}$ to force certain edge orientations in $\rho$. The incompatible orientations will consist of a binary out-directed tree $T_{0}$ and a binary out-directed tree with the successor-closures of two vertices identified, which we denote by $T_{1}$: these cannot start from the same point of a $2$-orientation, as one contains a $4$-cycle and the other does not. The idea to use two incompatible orientations in the WAP commutative diagram comes from the proof of [4, Theorem 5.2].
@@ 4.5. $(\mathcal{J},\leq_{1})$ does not have WAP
Suppose $(\mathcal{J},\leq_{1})$ has WAP, seeking a contradiction. Let $\{a_{0}\}\in\mathcal{J}$ be a singleton with the trivial linear order. Then there exists $\{a_{0}\}\leq_{1}A^{\prec}\in\mathcal{J}$ with $A^{\prec}$ witnessing WAP for $\{a_{0}\}$. Take $A\leq_{1}B\in\mathcal{C}_{1}$ witnessing for $A$ the expansion property of $\mathcal{J}$ over $(\mathcal{C}_{1},\leq_{1})$. (Here we use Theorem 2.30, recalling that $Y=X(\mathcal{J})$ is a minimal $G$-flow.)
@@ 4.6. $\mathcal{LO}(M_{1})$ is not minimal
We will show that $\mathcal{Q}_{1}$ is a reasonable class of expansions of $(\mathcal{C}_{1},\leq_{1})$ (see Definition 2.21). Parts (2) and (3) of reasonableness are immediate. For parts (1) and (4), take $A^{\prec}\in\mathcal{Q}_{1}$ and $B\in\mathcal{C}_{1}$ with $A\leq_{1}B$ (where we allow $A^{\prec}=\varnothing$). Let $\tau_{A}$ be the acyclic $2$-orientation induced by $\prec_{A}$ on $A$. As $A\leq_{1}B$, there exists an acyclic $2$-orientation $\tau_{B}$ of $B$ extending $\tau_{A}$. Let $\prec_{0}=\{(b,b^{\prime})\in B^{2}:b\neq b^{\prime}$ and there exists an out-path from $b^{\prime}$ to $b$ in $\tau_{B}\}$. Then $\prec_{0}$ is a strict partial order on $B$. $\prec_{A}$ and $\prec_{0}$ are compatible, and so we may extend the partial order $\prec_{A}\cup\prec_{0}$ arbitrarily to a linear order $\prec_{B}$ on $B$. Then $\prec_{B}$ induces $\tau_{B}$, so $(B,\prec_{B})\in\mathcal{Q}_{1}$.
By Lemma 2.26, we therefore have that $X(\mathcal{Q}_{1})$ is a subflow of $\mathcal{LO}(M_{1})$. To see that it is a proper subflow, we produce a linear order on $M_{1}$ which does not induce an acyclic $2$-orientation. Let $\prec$ be the linear order of the Fraïssé limit of the order expansion $(\mathcal{C}_{1}^{\prec},\leq_{1})$ of $(\mathcal{C}_{1},\leq_{1})$. By genericity, there exists a graph $A\leq_{1}M_{1}$ consisting of vertices $a,b_{1},\cdots,b_{3}$ and edges $ab_{i}$ with $b_{i}\prec a$ ($1\leq i\leq 3$), so $\prec$ does not induce a 2-orientation. ∎
See [18, Section 4.6] for an explicit example of a minimal subflow of $\mathcal{LO}(M_{1})$.

### v2

@@ 4.3. Proof idea - informal overview
The incompatibility of the two ordered graphs $C_{0}^{\prec},C_{1}^{\prec}$ in $\mathcal{J}$ will result from them forcing incompatible orientations: we can use the order $\prec_{\beta}$ to force certain edge orientations in $\rho$. The incompatible orientations will consist of a binary out-directed tree $T_{0}$ and a binary out-directed tree with the successor-closures of two vertices identified, which we denote by $T_{1}$: these cannot start from the same point of a $2$-orientation, as one contains a $4$-cycle and the other does not. The idea to use two incompatible orientations in the WAP commutative diagram comes from the proof of [5, Theorem 5.2].
@@ 4.5. $(\mathcal{J},\leq_{1})$ does not have WAP
Suppose $(\mathcal{J},\leq_{1})$ has WAP, seeking a contradiction. Let $\{a_{0}\}\in\mathcal{J}$ be a singleton with the trivial linear order. Then there exists $\{a_{0}\}\leq_{1}A^{\prec}\in\mathcal{J}$ with $A^{\prec}$ witnessing WAP for $\{a_{0}\}$. Take $A\leq_{1}B\in\mathcal{C}_{1}$ witnessing for $A$ the expansion property of $\mathcal{J}$ over $(\mathcal{C}_{1},\leq_{1})$. (Here we use 2.2, recalling that $Y=X(\mathcal{J})$ is a minimal $G$-flow.)
@@ 4.6. $\mathcal{LO}(M_{1})$ is not minimal
We will show that $\mathcal{Q}_{1}$ is a reasonable class of expansions of $(\mathcal{C}_{1},\leq_{1})$ (see [5, Definition 2.14]). Parts (2) and (3) of reasonableness are immediate. For parts (1) and (4), take $A^{\prec}\in\mathcal{Q}_{1}$ and $B\in\mathcal{C}_{1}$ with $A\leq_{1}B$ (where we allow $A^{\prec}=\varnothing$). Let $\tau_{A}$ be the acyclic $2$-orientation induced by $\prec_{A}$ on $A$. As $A\leq_{1}B$, there exists an acyclic $2$-orientation $\tau_{B}$ of $B$ extending $\tau_{A}$. Let $\prec_{0}=\{(b,b^{\prime})\in B^{2}:b\neq b^{\prime}$ and there exists an out-path from $b^{\prime}$ to $b$ in $\tau_{B}\}$. Then $\prec_{0}$ is a strict partial order on $B$. $\prec_{A}$ and $\prec_{0}$ are compatible, and so we may extend the partial order $\prec_{A}\cup\prec_{0}$ arbitrarily to a linear order $\prec_{B}$ on $B$. Then $\prec_{B}$ induces $\tau_{B}$, so $(B,\prec_{B})\in\mathcal{Q}_{1}$.
By [5, Theorem 2.15], we therefore have that $X(\mathcal{Q}_{1})$ is a subflow of $\mathcal{LO}(M_{1})$. To see that it is a proper subflow, we produce a linear order on $M_{1}$ which does not induce an acyclic $2$-orientation. Let $\prec$ be the linear order of the Fraïssé limit of the order expansion $(\mathcal{C}_{1}^{\prec},\leq_{1})$ of $(\mathcal{C}_{1},\leq_{1})$. By genericity, there exists a graph $A\leq_{1}M_{1}$ consisting of vertices $a,b_{1},\cdots,b_{3}$ and edges $ab_{i}$ with $b_{i}\prec a$ ($1\leq i\leq 3$), so $\prec$ does not induce a 2-orientation. ∎
See [16, Section 4.6] for an explicit example of a minimal subflow of $\mathcal{LO}(M_{1})$.